# bioSBM: a random graph model to integrate epigenomic data in chromatin structure prediction

### v1

@@ II.3 Posterior inference
where $Y$ and $X$ are for the OE Hi-C interaction data and the biochemical covariates respectively, $\eta_{1:N}$ are the latent membership vectors ($\theta_{1:N}$ are the normalized versions), $Z$ represents the community assignments for edges, $\Psi$ includes the global model parameters, and $\mathbb{E}_{q}[\cdot]$ denotes the expectation value with respect to the variational distribution $q$. The variational distribution $q(\eta_{1:N},Z)$ approximates the true posterior distribution $P(\eta_{1:N},Z|Y,X,\Psi)$.
The algorithm proceeds in two main steps:
@@
The iterative process of alternating between the E-step and the M-step continues until convergence, at which point the model parameters and variational distributions jointly provide an interpretation of the chromatin interaction patterns. For the detailed mathematical derivations and specific parameter update rules, refer to Sec. S1 in SM.
@@ III Results
To train the model, we perform posterior inference using $M$ pairs of biochemical covariates and Hi-C matrices $(X^{\mu},Y^{\mu})_{\mu=1}^{M}$, computing approximate posterior distributions over per-node latent membership vectors $\theta_{i}$ and the model parameters. In particular, the two sets of chromosomes we use are the odd-numbered chromosomes from 11 to 21 and the even-numbered chromosomes from 12 to 22. We use the model trained on one set to make predictions on the other and vice-versa. We applied the inference algorithm to data from the GM12878 lymphoblastoid cell line at a resolution of 100 kbps, and through Bayesian model selection, using evidence lower bound (ELBO) as a criterion, we determined that the optimal number of communities for our model is $K=12$. This suggests a more complex organization structure that extends beyond the conventional binary A/B compartmentalization and the more nuanced subdivision in subcompartments (up to six) [9]. At the same time, our model provides a more concise representation of chromatin contacts than the 30 epigenetically distinct “binding domains” suggested by Esposito et al [42] using polymer modeling.
@@
To evaluate the biological relevance of the communities identified by our model, we performed $k$-means clustering on the MAP membership vectors $\theta_{i}$. Clustering into two groups allowed us to compare these clusters to known A/B compartments, while varying the number of clusters, depending on the number of significantly present subcompartments in each chromosome, enabled comparisons with the subcompartments defined by Rao et al. [9].
Figure 2: The latent representation found through inference of the bioSBM model is biologically meaningful. The top row in each plot represents the clusters obtained by applying k-means on the MAP membership vectors inferred through our algorithm. The bottom rows are biological annotations.
Figure 3: Log fold enrichment in biochemical features (such as the presence of histone marks) for chromosome 19. The patterns of enrichment for our clusters are in good agreement with the enrichment found for A/B compartments and subcompartments.
The results of the clustering showed a significant overlap with known biological annotations. Fig. 2 illustrates this comparison for chromosomes 16 and 19. Both the binary subdivision in A/B compartments and the more granular classification in subcompartments can be captured from the full mixed membership vectors inferred by our model. Further validation of the inferred communities was performed by assessing the enrichment of each k-means-derived cluster in the biochemical features $x_{i}$. These enrichments were compared to those observed in A/B compartments and subcompartments annotations (see Fig. 3), where we can see a near-perfect agreement of bioSBM results with the enrichment of the binary A/B classification and a very good agreement with the enrichment of the subcompartments defined by [9].
@@
Figure 4: (a) Enrichment of mixed membership communities in the biochemical features used in the model. Each community correlates with distinct epigenetic features. (b) Interaction intensities between all pairs of communities. (c) Membership proportions of communities 4 and 5 as a function of node indices.
Importantly, the model goes beyond merely segmenting chromatin regions based on their epigenetic features; it also illustrates how these different classes interact, as represented by the matrix $B$, where each entry $B_{kg}$ encodes the interaction strength between class $k$ and $g$. Fig. 4(a) shows that each of the 12 inferred communities strongly correlates with distinct epigenetic marks. Interestingly, despite the apparent redundancy for some of the communities at the level of epigenomic profiles, there are clear distinctions between the interaction patterns of different classes. For instance (see Fig. 4(b)), classes 0 and 2 have similar epigenetic profiles but regions predominantly associated with class 0 have reduced contact frequency with regions primarily linked to class 2, and vice versa. Another notable observation is that classes 4 and 5 share a similar enrichment profile but differ in interaction patterns. Class 4 interacts only with itself and the epigenetically similar class 5, while class 5 interacts positively with classes 6, 7, and 8 as well. The inset of Fig. 4(c) shows that regions with non-zero probabilities of belonging to classes 4 and 5 overlap with clusters corresponding to the B2 and B4 compartments. Here however nodes need not be categorized into one class or another, they can share properties of multiple communities in different proportions.
Altogether these observations show that the bioSBM representation of chromatin interaction patterns provides a more nuanced description than the one provided by simple segmentation of the genome in different clusters.
III.2 bioSBM predicts the structure of the chromatin in unseen cell lines
@@
Specifically, given matrices $B$ and $\Gamma$, inferred from some training chromosomes in some conditions, and biochemical measurements for test chromosomes (either different chromosomes in the same cell line, or chromosomes from a different cell line), we can compute their community structure as $\theta_{i}=\Gamma x_{i}$. With the predicted community structure, we use the matrix $B$ to sample contact maps and compute their expectations, which can then be compared to experimentally obtained Hi-C maps (refer to Sec. S3 in SM for technical details).
Figure 5: The figure shows the Hi-C maps and the Pearson correlation matrix for chromosome 19. These maps are computed with the biochemical covariates associated with the chromosome and the parameters for the stochastic block model learned on a different set of chromosomes. The bottom triangle of each map is the experimentally observed Hi-C or Pearson correlation matrix, the top triangles are the model-generated counterparts. Above each map the colored bars represent the segmentation of the genome arising from k-mean clustering of the model mixed-membership vectors, while the bars on the left of each map are the subcompartment annotations from [9]
Fig. 5 illustrates the testing pipeline, providing an example of the predicted Hi-C map and predicted Pearson correlation matrix for chromosome 19. Here, biochemical covariates associated with chromosome 19 were used, while the block model interaction parameters were inferred from a different set of chromosomes. The strong correlation between experimental and predicted maps, as quantified by Pearson correlation coefficients (Fig. 6), demonstrates the model’s predictive accuracy. Interestingly, the difference in accuracy between the training and test sets was marginal, with a one-tailed Mann-Whitney U test yielding a p-value of 0.12. Therefore, bioSBM effectively generalizes across different chromosomes within the same cell line, suggesting that the inferred associations $B_{kg}$ between communities and the biochemistry-to-structure map $\Gamma$ reflect genuine chromatin interactions.
Further support for this conclusion comes from a comparison with state-of-the-art results reported by Esposito et al (2022) [42], where a code linking 1D chromatin marks to the interaction parameters of a polymer model was learned. While the polymer model similarly recapitulates chromatin contact patterns in training data, testing on unseen chromosomes from the GM12878 cell line strongly favors the bioSBM model in terms of predictive performance (see Fig. 6).
@@
Figure 6: Box plots of correlations between experimental maps and model generated maps, where the individual data points are single chromosomes. White diamonds represent the means of the box plots and the ivory squares are the correlation values for the example chromosome 11. The insets show comparisons of experimental maps (lower triangular part) and model-generated maps (upper triangle). To test whether the training performance was significantly higher or not than the test on different chromosomes on the same cell line we performed a one-tailed Mann-Whitney U test, that yielded a negative answer ($p$-value = 0.12). For tests on HCT116 we see that the correlations are higher in the RAD21 depleted scenario (tested for significance with a $p$-value = 0.03). The green box plots refer to correlations reported from Esposito et al. (2022) [42] for a polymer-based model on the same dataset.
@@
Finally, we tested the model on data from the human colorectal carcinoma cell line HCT116, where auxin-inducible degron (AID) technology was used to degrade RAD21 [43]. We apply the model as trained on GM12878 to independent covariates from the ENCODE Project [44] obtained from esperiments on the HCT116 cell line without AID-induced RAD21 degradation. In this case, we observe a decrease in predictive accuracy, likely due to the mismatch between the learned map $\Gamma$, which includes RAD21, and a biological sample where the latter has been degraded. Additionally, RAD21 deletion could have induced slight changes also in the other covariates, further affecting predictive performance. Notably, when the RAD21 covariate was excluded from the model (by setting the corresponding row in the feature matrix $X$ to zero), performance improved significantly ($p$-value = 0.03), suggesting the flexibility and robustness of bioSBM in capturing chromatin interactions under different cellular conditions.
@@ IV Discussion and conclusions
The bioSBM model introduced in this study offers a novel approach for modeling long-range chromatin interactions by integrating Hi-C data with biochemical covariates such as histone modifications and binding of transcription factors. By employing a mixed membership stochastic block model, we capture a more refined and nuanced view of chromatin structure, extending beyond the traditional binary A/B and sub-compartments framework. Our results show that the 12 communities identified by bioSBM correlate with known epigenetic features, reinforcing the idea that chromatin interactions are closely tied to the biochemical landscape of the genome. The partial agreement with subcompartments suggests that our model may capture additional layers of chromatin interaction complexity that may be missed by conventional classification methods.
@@
Although more abstract than polymer models, which explicitly take into account (to varying degrees of detail) the physical nature of the linear chromosomes, the latent representation learned by bioSBM remains biologically interpretable. The inferred associations between biochemistry and structure encoded in the linear map $\Gamma$ and the communities interactions contained in $B$ provide a starting point for systematically exploring more mechanistic descriptions of how chromatin folding is affected by nuclear biochemical processes.
@@
The predictive power of the model is another important contribution. By leveraging biochemical covariates, bioSBM can accurately predict chromatin contacts across different chromosomes and cell lines, comparing favorably with state-of-the-art polymer approaches [42]. Notably, the model’s robustness is highlighted by its performance on the HCT116 cell line, where the removal of the RAD21 covariate improved predictive accuracy, indicating that bioSBM can adapt to different chromatin environments and capture interactions under varying conditions, such as RAD21 depletion. Interestingly, bioSBM also compares favorably with recent deep-learning models. Many deep learning approaches predict Hi-C maps using only DNA sequence data as input features [14, 15], which limits their ability to capture cell-type specific variations. In contrast, bioSBM explicitly incorporates epigenetic features, making it more suited for cell-type-specific predictions. Additionally, bioSBM’s predictive power is comparable to a recent deep learning model that does include epigenetic signals [17]. However, deep learning models have a distinct advantage in terms of computational power, allowing them to predict Hi-C maps at higher resolutions and model small-scale perturbations, such as the effects of single nucleotide polymorphisms (SNPs) [14]. While our model cannot operate at this fine-grained scale, it remains effective in predicting the effects of larger-scale manipulations such as the AID-induced RAD21 reported above. Crucially, bioSBM stands out for its interpretability: every parameter has a clear probabilistic semantic, allowing their analysis to uncover a clear biological meaning. This makes bioSBM a valuable, flexible tool for studying chromatin organization in diverse cellular contexts.
@@
While this study focuses on Hi-C data, the so-called enrichment methods [45], such as Promoter Capture Hi-C (PCHi-C) [46] or ChIA-PET [47, 48], may offer more functionally relevant perspectives on chromatin structure. PCHi-C, for example, highlights promoter regions interactions, while ChIA-PET focuses on interactions involving specific proteins. These methods present challenges for polymer models because they capture interactions between non-contiguous regions. However, bioSBM, being graph-based, could accommodate these non-contiguous interactions with some adaptation. Though some graph-based studies exist that link nuclear biochemistry to chromatin structure as assayed by these enrichment-based data [21, 22], a fully generative predictive model mapping biochemistry to the chromatin interaction patterns, such as that provided by bioSBM for Hi-C, has yet to be developed. With appropriate adjustments, bioSBM could readily be extended to accommodate these data types.

### v2

@@ II.3 Posterior inference
where $Y$ and $X$ are for the OE Hi-C interaction data and the biochemical covariates respectively, $\eta_{1:N}$ are the latent membership vectors ($\theta_{1:N}$ are the normalized versions), $Z$ represents the community assignments for edges, $\Psi$ includes the global model parameters, and the symbol $\mathbb{E}_{q}(\cdot)$ denotes the expectation value of the bracketed quantity with respect to the variational distribution $q$. The variational distribution $q(\eta_{1:N},Z)$ approximates the true posterior distribution $P(\eta_{1:N},Z|Y,X,\Psi)$.
Then, the algorithm proceeds in two main steps:
@@
The iterative process of alternating between the E-step and the M-step continues until convergence, at which point the model parameters and variational distributions jointly provide an interpretation of the chromatin interaction patterns. For the detailed mathematical derivations and specific parameter update rules, refer to Sec. S1 in SM [35].
@@ III Results
To train the model, we have performed posterior inference using pairs of biochemical covariates and Hi-C matrices $(X^{\mu},Y^{\mu})_{\mu=1}^{M}$ for $M$ chromosomes ($M=11$ was chosen as the number of chromosomes in each training set in the experiments). More specifically, we employ the two sets of odd-numbered human chromosomes from 1 to 21 and the even-numbered human chromosomes from 2 to 22, and we use the model trained on one set to make predictions on the other and vice versa computing approximate posterior distributions over per-node latent membership vectors $\theta_{i}$ and the model parameters. We have applied the inference algorithm to data from the GM12878 lymphoblastoid cell line at a resolution of 100 kilo-basepairs; then, through Bayesian model selection, using evidence lower bound (ELBO) as a criterion, we determined that the optimal number of communities for our model is $K=7$ (see Fig. S1 in SM [35]). These results align with recent efforts to describe chromatin organization extending beyond the conventional A/B compartmentalization [6] and encompassing more nuanced frameworks, including subcompartments [9] and Interaction Profile Groups (IPGs) from Spracklin et al. [15]. Notably, a recent orthogonal approach based on polymer modeling by Esposito et al. [54] showed that their model could recapitulate Hi-C contact patterns with a set of “binding domains”, which could be clustered into 9 statistically significant, epigenetically distinct groups. The similarity in the number of inferred domain types, despite the methodological differences, highlights the consistency and biological relevance of both approaches.
@@
Figure 2: The latent representation found through inference of the bioSBM model is biologically meaningful. The top rows in (a) and (b) represent the clusters obtained by applying $k$-means on the MAP membership vectors inferred through our algorithm. The bottom rows are biological annotations. (c) Adjusted Mutual Information score between the clustering obtained by bioSBM’s membership vectors and annotations previously reported in the literature, for the example chromosomes 16 and 19 and for all other chromosomes from 1 to 22.
Figure 3: Log-fold enrichment in biochemical features (such as the presence of histone marks) for chromosome 19. The patterns of enrichment for our clusters are in good agreement with the enrichment found for A/B compartments, subcompartments, and IPGs from [15]. The naming for the cluster labels for the clustering obtained applying $k$-means to bioSBM membership vectors, and for IPGs, have been mapped to match the nomenclature of subcompartments by [9].
@@
The first output of the inference procedure is a set of $K$-dimensional ($K=7$ for GM12878) mixed-membership probability vectors $\theta_{i}$ for all the genomic bins in the training data. These vectors represent the probabilistic community membership of each genomic region.
To facilitate biological interpretation, we perform a separate clustering step on these vectors using $k$-means clustering [55], to obtain discrete labels. Clustering into two groups allowed us to compare these clusters to known A/B compartments, while setting the number of clusters to 6 enabled comparisons both with the subcompartments defined by Rao et al. [9] and the Interaction Profile Groups (IPGs) computed with the algorithm by Spracklin et al. [15] (see Sec. S2.E in SM [35]).
The results of the clustering showed a significant overlap with the established biological annotations. Fig. 2 illustrates this comparison for chromosomes 16 and 19. The binary subdivision in A/B compartments as well as the more granular classification in subcompartments or IPGs can be captured from the full mixed membership vectors inferred by our model (see Fig. S9 in SM [35] a comparison of centrome distances of the various clusters.). Further validation of the inferred communities was performed by assessing the enrichment (see Sec. S2.C in SM [35] for details) of each $k$-mean-derived cluster in the biochemical features $x_{i}$. These enrichments were compared (see Fig. 3) to those observed in A/B compartments, subcompartments, and annotations from [15], where we can see a near-perfect agreement of the discretized bioSBM results with the enrichment of the binary A/B classification and a very good agreement with the enrichment of the subcompartments defined by [9]and the IPGs..
@@
Figure 4: (a) Enrichment of mixed membership communities in the biochemical features used in the model. Each community correlates with distinct epigenetic features. (b) Interaction intensities between all pairs of communities. (c) Example of membership proportions of communities 4 and 5, for whole chromosome 19.
Importantly, the model goes beyond merely segmenting chromatin regions based on their epigenetic features; it also illustrates how these different communities interact, as represented by the matrix $B$ where each entry $B_{kg}$ encodes the interaction strength between community $k$ and $g$. Panel (a) in Fig. 4 shows that each of the $7$ inferred communities is associated both with distinct epigenomic patterns and with interaction patterns between communities. Interestingly, despite the apparent redundancy for some of the communities at the level of epigenomic profiles, there are clear distinctions between the interaction patterns of different communities. For instance (see panel (b) in Fig. 4), communities 0 and 1 have similar epigenetic profiles but regions predominantly associated with community 0 interact mainly with community 0 and community 1, while nodes predominantly associated to community 1 interact also with community 5. Another notable observation is that community 2 interacts positively only with the epigenetically similar communities 3 and 4, but community 3 interacts positively also with the very different community 5. The inset of panel (c) in Fig. 4 shows that regions with non-zero probabilities of belonging to communities 2 and 3 overlap with clusters corresponding to the B2 and B4 compartments. Here, however, nodes do not need to be categorized into one community or another, as they can share properties of multiple communities in different proportions (see also Fig. S2 in SM[35] for some results characterizing the importance and degree of the nodes’ mixed-membershipness).
Altogether, these observations show that the bioSBM representation of chromatin interaction patterns provides a more nuanced description than the one provided by simple segmentation of the genome in different clusters.
III.2 bioSBM’s predictive power
@@
Specifically, given matrices $B$ and $\Gamma$, inferred from some training chromosomes in some conditions, and biochemical measurements for test chromosomes (either different chromosomes from the same cell line, or chromosomes from a different cell line), we can compute their community structure as $\theta_{i}=\Gamma x_{i}$. With the predicted community structure, we use the matrix $B$ to sample contact maps and compute their expectations, which can then be compared to experimentally obtained Hi-C maps (refer to Sec. S2.B in SM [35] for technical details and panels (b) and (c) in Fig. 1 for a schematic representation of the pipeline).
III.2.1 bioSBM assessment strategy
Testing machine learning models in biology is non-trivial, because the meaning of generalisation is often unclear (see e.g. Schreider et al. [56] for an in-depth discussion): one may seek to predict for unseen regions in the same chromosome (hence, with similar local environment), or for different chromosomes in the same biological condition (cell line, tissue, etc.), or seek to extrapolate to completely new unseen conditions. In our experiments, we stress-test bioSBM under the two most stringent conditions: first, we learn model parameters using ChIP-seq and Hi-C data from a subset of chromosomes in the lymphoblastoid cell line GM12878. Then, we test model performance in predicting Hi-C data on the remaining chromosomes, using as input the relative ChIP-seq data (panels (b) and (c) in Fig. 1). We then seek to assess the performance of bioSBM on the task of predicting Hi-C data in two completely unseen cell lines: the K562 leukemia cell line, using ENCODE ChIP-seq data, and the colorectal carcinoma line HCT116, with and without a major external perturbation (the rapid degradation of the RAD21 loop extrusion factor [57]). On the HCT116 RAD21-cell line, we initially perform the prediction task using the parameters learned from the GM12878 data set and ENCODE ChIP-seq data from HCT116, including tracks for RAD21. We then seek to simulate the ability of bioSBM to predict perturbation by repeating our prediction experiment (still with the same parameters), but simply setting to zero the ChIP-seq tracks for RAD21 (see panels (b) and (c) in Fig. 5). This enables us to assess bioSBM predictions both under covariate shift (moving from GM12878 to HCT116 RAD21-cells), and its ability to model perturbations by changing the inputs in a deterministic way (setting to zero the RAD21 tracks).
III.2.2 bioSBM predicts the structure of the chromatin in unseen cell lines
The results of the tests described above are reported in panel (a) in Fig. 5. The first 2 columns show Pearson correlation between experimental log O/E Hi-C maps, and model-generated maps, for the training chromosomes, and for test chromosomes on the same cell line. Interestingly, the difference in accuracy between the training and test sets was marginal, with a one-tailed Mann-Whitney U test yielding a $p$-value of 0.82 (meaning the difference in accuracy is not statistically significant), see Fig. S5 in SM [35] for a comparison between predicted and real maps for all chromosomes. Therefore, bioSBM effectively generalizes across different chromosomes within the same cell line, suggesting that the inferred associations $B_{kg}$ between communities and the biochemistry-to-structure map $\Gamma$ reflect genuine chromatin interactions. Despite some differences in methodology and evaluation metrics, the correlation values we observe align well with the “distance-corrected” Pearson correlation values reported by Esposito et al. [54]. Their polymer-based approach, grounded in physical modeling of chromatin structure, offers valuable and direct mechanistic insight. While their model performs well on training data, it exhibits a noticeable decrease in predictive accuracy on unseen chromosomes. In contrast, bioSBM maintains robust predictive performance across chromosomes, highlighting its strong generalization capability. These complementary strengths illustrate the potential for integrating diverse modeling strategies to better understand chromatin organization.
@@
Figure 5: (a) Box plots of correlations between experimental maps and model generated maps, where the individual data points are single chromosomes. Ivory squares are the correlation values for an example chromosome 11. To test whether the training performance was significantly higher or not than the test on different chromosomes on the same cell line we performed a one-tailed Mann-Whitney U test, that yielded a negative answer ($p$-value = 0.82). To test whether the improvement of predictions granted by removal of the RAD21 feature, for predictions on the AID-treated HCT116 cell line, was significant, we used the same test, which yielded a $p\text{-value}=1.20\cdot 10^{-4}$ (significant difference). Also the predictive accuracy on HCT116 without AID treatment is significantly higher ($p\text{-value}=2.33\cdot 10^{-5}$). The orange box plots refer to distance-corrected correlations reported from Esposito et al. [54] for a polymer-based model on the same dataset, see also Fig. S6 in SM [35] for a comparison with the results obtained with a neural network parametrization of $\Gamma$. (b) and (c) Schematization of the test performed on the AID-treated HCT116 cell line.
Additionally, we apply the model trained on half the chromosomes of GM12878, to make predictions on contacts on the other half of the chromosomes of the cancer cell lines HCT116 and K562 (last and third to last columns in panel (a) of Fig. 5), from covariate tracks downloaded from the ENCODE [58] database. The results show that the model trained on GM12878 chromosomes is able to explain a large fraction of the variance in the data from HCT116 and K562. Finally, we tested the model on a dataset for the HCT116 cell line, where auxin-inducible degron (AID) technology was used to degrade RAD21 [57].
The tests on HCT116 are especially informative, as Rao et al. [57] showed that RAD21 depletion in HCT116 cells preserves the global compartmentalization pattern–quantified in terms of the first eigenvector of the Hi-C correlation maps–but introduces quantitative changes, such as a marked increase in compartment strength. To quantify this effect, we computed chromosome-wise correlations between treated and untreated Hi-C maps for the HCT116 cell line (data from Rao et al. 2017) [57], yielding an average Pearson correlation of $\sim 0.78$. This confirms that while the plaid pattern remains globally consistent, substantial quantitative differences exist that a predictive model like bioSBM should aim to capture.
We make predictions of contact frequencies with the model trained on GM12878 and independent covariates from the ENCODE Project [58] obtained from experiments on the HCT116 cell line (without AID-induced RAD21 degradation). We tested two scenarios: one where all covariates were used (denoted “RAD21 yes”, and one where the RAD21 track was zeroed out to simulate degradation (“RAD21 no”). We then compare these predicted maps to experimental maps of HCT116 cells without AID treatment (and therefore still have RAD21) or cells after six hours of treatment, leading to RAD21 depletion. We denote the first experimental dataset as RAD21+, and the second one as RAD21- (see Table S1 in SM [35], for data accession codes).
@@
In the RAD21 yes vs RAD21+ setup, we observe a decrease in predictive accuracy, compared to predictions made on the same cell line as the training data, likely due to the mismatch between the learned map $\Gamma$, which includes RAD21, and a biological sample where the latter has been degraded. Additionally, RAD21 deletion could have induced slight changes also in the other covariates, further affecting predictive performance. Notably, when the RAD21 covariate was excluded from the input (“RAD21 no”), performance on the RAD21- experimental data improved significantly ($p$-value = $1.20\cdot 10^{-4}$) (see Fig. S8 in SM [35] for the analysis stratified by genomic distance). Conversely, the in silico RAD21 exclusion, reduced the predictive accuracy for the RAD21+ data (see Fig. 5).
These results underscore the flexibility and robustness of bioSBM in capturing chromatin interactions under different cellular conditions (see also Fig. S7 in SM [35] for an example of predictions with a model retrained on a subset of covariates).
@@ IV Discussion and conclusions
The bioSBM model introduced in this study offers a novel approach for modeling long-range chromatin interactions by integrating Hi-C data with biochemical covariates such as histone modifications and binding of transcription factors. By employing a mixed membership stochastic block model, we capture a more refined and nuanced view of chromatin structure, extending beyond the traditional binary A/B and sub-compartments framework. Our results show that the 7 communities identified by bioSBM correlate with known epigenetic features, reinforcing the idea that chromatin interactions are closely tied to the biochemical landscape of the genome. The partial agreement with subcompartments suggests that our model may capture additional layers of chromatin interaction complexity that may be missed by conventional classification methods.
@@
Although more abstract than polymer models, which explicitly take into account (to varying degrees of detail) the physical nature of the linear chromosomes, the latent representation learned by bioSBM remains biologically interpretable. The inferred associations between biochemistry and structure encoded in the linear map $\Gamma$ and the communities interactions contained in $B$ provide a starting point for systematically exploring more mechanistic descriptions of how chromatin folding is affected by nuclear biochemical processes. Importantly, bioSBM does not incorporate the underlying polymeric nature of chromatin, which is known to influence 3D genome architecture. Factors such as the dynamics of loop extrusion, maintained by SMC complexes, and the local mechanical state of the chromatin fiber can lead to distinct contact patterns that are not always explained by epigenomic marks alone. This has been demonstrated both experimentally (e.g., [59]) and via polymer simulations (e.g., [60]). These properties – absent from our current modeling framework – may account for some of the structural distinctions inferred by bioSBM, even among communities with seemingly similar epigenomic profiles.
@@
The predictive power of the model is another important contribution. By leveraging biochemical covariates, bioSBM can accurately predict chromatin contacts across different chromosomes and cell lines, comparing favorably with state-of-the-art polymer approaches [54]. Notably, the model’s robustness is highlighted by its performance on the HCT116 RAD21-cell line, where the removal of the RAD21 input covariate improved predictive accuracy, indicating that bioSBM can adapt to different chromatin environments and capture interactions under varying conditions, such as RAD21 depletion. Interestingly, bioSBM also compares favorably with recent deep-learning models, though their objectives and designs differ. While some deep learning approaches rely solely on DNA sequence input [20, 21], limiting their ability to model cell-type specific variations, bioSBM explicitly integrates epigenetic features for this purpose. A recent model, Epiphany [23], also incorporated epigenomic data and reports predictive performance on held-out chromosomes in GM12878 that is in a similar range to what we observe for bioSBM. However, we note that a direct comparison is not straighforward, as Epiphany operates at a higher resolution and focuses on fine-scale perturbations. Additionally, its evaluation in other cell types is based on structural features such as insulation scores rather than whole-map correlation.
While our model cannot operate at this fine-grained scale, it remains effective in predicting the effects of larger-scale manipulations such as the AID-induced RAD21 reported above. Despite these differences, both approaches highlight the value of integrating epigenomic information into 3D genome prediction models. Crucially, bioSBM stands out for its interpretability: every parameter has a clear probabilistic semantic, allowing their analysis to uncover a clear biological meaning. This makes bioSBM a valuable, flexible tool for studying chromatin organization in diverse cellular contexts.
@@
While this study focuses on Hi-C data, the so-called enrichment methods [61], such as Promoter Capture Hi-C (PCHi-C) [62] or ChIA-PET [63, 64], may offer more functionally relevant perspectives on chromatin structure. PCHi-C, for example, highlights promoter region interactions, while ChIA-PET focuses on interactions involving specific proteins. These methods present challenges for polymer models because they capture interactions between non-contiguous regions. However, bioSBM, being graph-based, could accommodate these non-contiguous interactions with some adaptation. Though some graph-based studies exist that link nuclear biochemistry to chromatin structure as assayed by these enrichment-based data [29, 30], a fully generative predictive model mapping biochemistry to the chromatin interaction patterns, such as that provided by bioSBM for Hi-C, has yet to be developed. With appropriate adjustments, bioSBM could readily be extended to accommodate these data types.
@@
V Code availability
An implementation of the code for the inference algorithm described in this paper is available at https://github.com/alex-chenyi-zhang/bioSBM-code.
Acknowledgements.
GS acknowledges co-funding from Next Generation EU, in the context of the National Recovery and Resilience Plan, Investment PE1 - Project FAIR “Future Artificial Intelligence Research”. This resource was co-financed by the Next Generation EU [DM 1555 del 11.10.22]. AR acknowledges financial support from PNRR Grant CN 00000013 CN-HPC, M4C2I1.4, spoke 7, funded by Next Generation EU. AG acknowledges financial support by MIUR PRIN-COFIN2022 grant 2022JWAF7Y.